mr.may traveled from home to New York by a car with 15 mph speed. While he was driving, his girlfriend called him to go back home. He suddenly drove back home with 25 mph speed. If this round-trip took 32 minutes. What is the average speed he drove for the round-trip?
Answer by stanbon(75887)

mr.may traveled from home to New York by a car with 15 mph speed. While he was driving, his girlfriend called him to go back home. He suddenly drove back home with 25 mph speed. If this round-trip took 32 minutes. What is the average speed he drove for the round-trip?
-----------------
Going DATA:
rate = 15 mph ; distance = x miles ; time = x/15 hrs
----------------------
Returning DATA:
rate = 25 mph ; distance = x miles ; time = x/25 hrs
----------------
Equation:
time + time = 32/60 hrs
x/15 + x/25 = 32/60
x/15 + x/25 = 16/30
Multiply thru by 150 to get:
10x + 6x = 80
15x = 80
x = 5 miles
----------------------
Average speed = total distance / total time
Ave speed = 10 miles/(32/60)hr = 600/32 = 18.75 mph
=======================================================
Cheers,
Stan H.